# Collapsing regular Riemannian foliations with flat leaves

### v1

@@ Abstract.
We present how to collapse a manifold equipped with a closed flat regular Riemannian foliation with leaves of positive dimension on a compact manifold, while keeping the sectional curvature uniformly bounded from above and below. From this deformation, we show that a closed flat regular Riemannian foliation with leaves of positive dimension on a compact simply-connected manifold is given by torus actions. This gives a geometric characterization of aspherical regular Riemannian foliations given by torus actions.
@@
Collapse, F-structures, A-foliations
@@
$*{}^{*}$his work was supported by a UKRI Future Leaders Fellowship [grant number MR/W01176X/1; PI: J Harvey], by the DFG (281869850, RTG 2229 “Asymptotic Invariants and Limits of Groups and Spaces”), a DGAPA postdoctoral Scholarship of the Institute of Mathematics of UNAM, and DFG-Eigenestelle Fellowship CO 2359/1-1
@@ 1. Introduction
In the setting of singular Riemannian foliations, Galaz-García and Radeschi in [22] introduced an analogous concept to torus actions via isometries by considering singular Riemannian foliations whose leaves are closed and aspherical manifolds. These foliations are called $A$-foliations and examples of such foliations include effective torus actions by isometries. In general due to [22] and [8], on a simply-connected manifold the leaves of an $A$-foliation are finitely covered by tori, and moreover, the leaves of highest dimension are all homeomorphic to a torus. Thus in this way, $A$-foliations resemble torus actions. But in [16] examples are given of $A$-foliations which are not given by torus actions. From this discussion, it follows that it is natural to ask if we may characterize those $A$-foliations given by a torus action.
@@
þ Consider $(M,\mathcal{F},g)$ a regular $B$-foliation on a compact simply-connected Riemannian manifold. Then the foliation $\mathcal{F}$ is induced by an isometric torus action.
To prove þA we compare regular $B$-foliations to the so called $\mathsf{F}$-structures and $\mathsf{N}$-structures introduced by Cheeger, Gromov and Fukaya in [6] and [5]. These are other types of foliations, and were introduced when studying the notion of “collapse with bounded curvature” onto metric spaces. We recall that an $\mathsf{F}$-structure is given by an open cover $\{U_{i}\}_{i\in\mathfrak{I}}$ of $M$, together with a $T^{k_{i}}$-action on a finite normal (or Galois) cover $\pi_{i}\colon\tilde{U}_{i}\to U_{i}$ for each index, such that the torus actions satisfy certain compatibility conditions when passing from one neighborhood to another, and they are compatible with the action of the deck-transformation group of $\pi_{i}$ (see for example [20, Definition 19.2]). An $\mathsf{F}$-structure induces a partition of $M$ into submanifolds called the orbits of the $\mathsf{F}$-structure. We say that a sequence of Riemannian metrics $\{g_{n}\}_{n\in\mathbb{N}}$ on a compact manifold $M$ collapses with bounded curvature if there exist constants $\lambda\leq\Lambda\in\mathbb{R}$ such that
@@
and for any $p\in M$ the injectivity radius of $g_{n}$ at $p$ goes to 00 as $n\to\infty$ (see for example [38, p. 69]). By [6] and [7] we have that $\mathsf{F}$-structures characterize collapse with bounded curvature: A compact manifold collapses with bounded curvature if and only if it admits an $\mathsf{F}$-structure whose orbits have positive dimension.
For the proof of þA, we prove that given a regular $B$-foliation $(M,\mathcal{F},g)$ on a compact manifold with leaves of positive dimension we may always collapse $M$ with bounded curvature and bounded diameter.
@@
þ Consider $(M,\mathcal{F},g)$ a regular non-trivial $B$-foliation on a compact Riemannian manifold. Then by shrinking the directions tangent to the leaves of $\mathcal{F}$ we may collapse with bounded curvature and uniformly bounded diameter the manifold $(M,g)$.
We recall that in the study of collapse with bounded curvature in [5], given such a collapsing sequence the authors construct $\mathsf{N}$-structures and highly symmetric $\mathsf{N}$-invariant Riemannian metrics which are bounded by the collapsing sequence (see þ2.17). When the manifold $M$ has finite fundamental group an $\mathsf{N}$-structure is given by an $\mathsf{F}$-structure (see þ2.14). In the case when $M$ has finite fundamental group we use the relationship between $\mathsf{N}$-structures and $\mathsf{F}$-structures, the invariant metrics and the explicit collapsing sequence given in the proof of þB to show that the geometry of $(M,\mathcal{F},g)$ where $\mathcal{F}$ is a $B$-foliation with leaves of positive dimension is approximated by geometries invariant under $\mathsf{F}$-structures. We can also point to a relationship between the tangent spaces of the leaves of the foliations and the tangent spaces of the orbits of the $\mathsf{F}$-structure.
@@
þ Consider $(M,\mathcal{F},g)$ a regular non trivial $B$-foliation on a compact $m$-dimensional Riemannian manifold with finite fundamental group and $m>2$. Then there exists a sequence of Riemannian metrics $\{\bar{g}^{\varepsilon}_{N}(\varepsilon)\}$ with each metric invariant under an $\mathsf{F}$-structure, and such that as $\varepsilon\to 0$ we have that $(M,\bar{g}^{\varepsilon}_{N}(\varepsilon))$ converges in the Gromov-Hausdorff sense to $(M,g)$. Moreover, given $p\in M$ fixed and taking $x\in T_{e}T^{k}$ fixed in the Lie algebra of the torus tracing the leaves of the $\mathsf{F}$-structure around $p$, we have that the action field $X^{\ast}(p)$ converges to a direction tangent to the leaf $L_{p}$.
We remark that by þB, given a regular $B$-foliation on a compact manifold $M$ with leaves of positive dimension we have a continuous family $g_{\delta}$ of Riemannian metrics collapsing with bounded curvature to the leaf space. Combining this with [40, Theorem 0.4] we conclude that we have an upper bound for the minimum of the sectional curvature of a compact manifold equipped with a $B$-foliation whose leaves have positive dimension.
@@
Let $(M,\mathcal{F},g)$ be a regular $B$-foliation on a compact manifold with leaves of positive dimension. Then we have for $\lambda$ equal to the infimum of the sectional curvature of $g$, that $\lambda\leq 0$.
@@
þA sheds some light on this question: Regular flat $A$-foliations (i.e. $B$-foliations) on simply-connected manifolds are given by torus actions, and thus are $\mathsf{F}$-structures.
We highlight that on a simply-connected manifold, an $\mathsf{F}$-structure is induced by a torus action (see for example [42, Proposition 3.1, Lemma 3.2 and Lemma 4.1]). Thus it is a pure $\mathsf{F}$-structure, and therefore a $B$-foliation. Still such $B$-foliations may have singular leaves. For example torus actions of cohomogeneity $2$ can have singular leaves [37], [8]. Nonetheless, when $M$ has finite fundamental group a regular $A$-foliation with flat leaves is the limit of $\mathsf{F}$-structures by þC.
We point out that not all regular $A$-foliations admit a Riemannian metric making it a $B$-foliation. In [16] the authors construct manifolds of dimension $m\geq 9$ with finite fundamental group and each equipped with a regular $A$-foliation whose leaves are homeomorphic to the $(m-4)$-dimensional torus but not diffeomorphic to the $(m-4)$-dimensional torus. These foliations do not admit any Riemannian metric making them into a $B$-foliation, since by [3] a flat manifold is homeomorphic to the torus if and only if it is diffeomorphic to the torus.
@@
þ Is there an $A$-foliation on a simply-connected manifold which is not given by a torus action?
@@
In [8, Theorem A] a negative answer to þ2 is given for $A$-foliations of codimension $2$. Namely in [8], the author explored the problem of whether we can compare the diffeomorphism type of two compact simply-connected manifolds with $A$-foliations via their leaf spaces equipped with local information of the foliation, focusing on the case of $A$-foliations of codimension $2$. The positive answer given in [8, Theorem C and Lemma 5.11] to this comparison problem together with the classification of tours actions of cohomogeneity $2$ in [37], allows us to conclude that an $A$-foliation of codimension $2$ on a compact simply-connected manifold is given up to a foliated homeomorphism by a torus action. Then, by analyzing the local information of the foliation we are able to improve the conclusion to a foliated diffeomorphism.
@@
A crucial point in the proof of [8, Theorem A] is that for an $A$-foliation of codimension $2$ on a compact simply-connected manifold the stratification of the leaf space is simple and can be explicitly described. In general, even for singular Riemannian foliations on vector spaces with finite dimension, it is not clear how to compare the foliations via their leaf spaces (see [23, Question 1.2]). We point out that the proof presented in [8] is of a topological nature and only uses the fact that the presence of a foliated Riemannian metric guarantees a nice local structure of the foliation on tubular neighborhoods of leaves (see [33]). Moreover, to be able extend the ideas in [8] to higher dimensions, we would require to impose extra topological conditions on $M$ and on the leaf space of the foliation. Thus, to the best of our knowledge, the approach presented in this manuscript is a different approach for comparing foliated structures to what exists in the literature.
@@
The present work was motivated by the need to give geometric conditions on $A$-foliations that force them to be given by torus actions. Due to [5] and [42] it follows that on simply-connected manifolds, torus actions with orbits of positive dimension are characterized by the phenomenon of collapse with bounded curvature (see þ2.17). Thus, a natural approach to finding sufficient geometric conditions for an $A$-foliation to be homogeneous, is to find sufficient geometric conditions which lets us collapse with bounded curvature the manifold by shrinking the leaves of a given $A$-foliation on a simply-connected manifold.
@@
We highlight that to the best of our knowledge in the literature there are few references on the subject of deforming foliations. In [14] the authors show that for a particular class of $1$-dimensional regular foliations, a portion of the manifold has a cover into “long and thin” submanifolds extending the notions in [5] to the foliated setting. In [9] the authors consider a parametrized family of Lie groupoids and give rigidity results for such “deformations” of Lie groupoids. Lie groupoids are closely related to regular foliations (see [34]), and also singular Riemannian foliations (see [2]). Applying these rigidity results for Lie groupoids, the authors in [11] consider on a compact manifold a family of regular foliations $\mathcal{F}_{t}$ parametrized by a real parameter $t\in[0,1]$, each having compact leaves and Hausdorff leaf space, and show that at $t=0$ the existence of a leaf $L$ with $H^{1}(L,\mathbb{R})=0$ is a sufficient condition for such a parametrization to be rigid, i.e. $\mathcal{F}_{t}=\mathcal{F}_{0}$. This is an analogous result to the conclusions presented in [13] and in [28]: Consider two regular foliations $\mathcal{F}_{1}$, and $\mathcal{F}_{2}$ with compact leaves on a fixed manifold $M$ without boundary, such that they are close to each other in an appropriate sense, and such that the leaf spaces of the foliations are Hausdorff spaces. In the case that there exists a leaf $L$ in one of the foliations with $H^{1}(L,\mathbb{R})=0$, there exists a foliated diffeomorphism between the foliated manifolds. For group actions there is a general rigidity result by Grove and Karcher [25]: Given a compact Lie group $G$, a compact manifold $M$, and two group actions $\mu_{1}\colon G\to\mathrm{Diff}(M)$ and $\mu_{2}\colon G\to\mathrm{Diff}(M)$, if the images $\mu_{1}(G)$ and $\mu_{2}(G)$ are $C^{1}$ close enough in $\mathrm{Diff}(M)$, then the images are conjugated in $\mathrm{Diff}(M)$.
@@
We point out that the proof of þA is also a rigidity result for a deformation procedure, since the proof consist of approximating our foliation by torus actions using þC, and showing that there exists a limit torus action whose orbits agree with the leaves of the foliation. Nonetheless, we point out that to the best of our knowledge there are no general rigidity results for singular Riemannian foliations as those presented in [11], [13], and [28]. Moreover, the rigidity result presented in þA agree with the conclusions in [25]. One can also compare þA with [40, Proposition 0.8] and comments after. In our work we also point to the existence of a fixed torus action on $M$, but we do show that we can take our foliated metric as the limit of the perturbed collapsing sequence.
@@
Our article is organized as follows: In Section 2 we review the necessary preliminaries on singular Riemannian foliations, as well as the concepts of $A$-foliations, $B$-foliations, and $\mathsf{N}$-structures. We also review Alexandrov spaces and equivariant convergence of metric spaces. At the end of Section 2, we present the relevant theorems of collapse theory used in our proofs. In Section 3 we give the proofs of our main theorems. We begin by presenting the proof of þB, and then proceed with the proof of þC. We use then these results and their proofs to prove þA.

### v2

@@ Abstract.
In this manuscript we present how to collapse a manifold equipped with a closed flat regular Riemannian foliation with leaves of positive dimension, while keeping the sectional curvature uniformly bounded from above and below. From this deformation, we show that in the case when the manifold is compact and simply connected the foliation is given by torus actions. This gives a geometric characterization of aspherical regular Riemannian foliations given by torus actions.
@@
Collapse, F-structures, A-foliations, B-foliations, singular Riemannian foliations
@@
This work was supported by a UKRI Future Leaders Fellowship [grant number MR/W01176X/1; PI: J Harvey], by the DFG (281869850, RTG 2229 “Asymptotic Invariants and Limits of Groups and Spaces”), a DGAPA postdoctoral Scholarship of the Institute of Mathematics of UNAM, and DFG-Eigenestelle Fellowship CO 2359/1-1 as part of the DFG-SPP 2026 “Geometry at Infinity”.
@@ 1. Introduction
In the setting of singular Riemannian foliations, Galaz-García and Radeschi in [22] introduced an analogous concept to torus actions via isometries by considering singular Riemannian foliations whose leaves are closed and aspherical manifolds. These foliations are called $A$-foliations and examples of such foliations include effective torus actions by isometries. In general due to [22] and [8], on a simply-connected manifold the leaves of an $A$-foliation are finitely covered by tori, and moreover, the leaves of highest dimension are all homeomorphic to a torus. Thus in this way, $A$-foliations resemble torus actions. But in [16] examples are given of $A$-foliations, on closed manifolds with a non-trivial finite fundamental group, which are not given by torus actions. From this discussion, it follows that it is natural to ask if we may characterize those $A$-foliations given by a torus action.
@@
Consider $(M,\mathcal{F},g)$ a regular $B$-foliation on a compact simply-connected Riemannian manifold. Then the foliation $\mathcal{F}$ is induced by an isometric torus action.
To prove ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions we compare regular $B$-foliations to the so called $\mathsf{F}$-structures and $\mathsf{N}$-structures introduced by Cheeger, Gromov and Fukaya in [6] and [5]. These are other types of foliations, and were introduced when studying the notion of “collapse with bounded curvature” onto metric spaces. We present now the definition of $\mathsf{F}$-structures (for the definition of $\mathsf{N}$-structures see Section 2.7). We recall that an $\mathsf{F}$-structure is given by an open cover $\{U_{i}\}_{i\in\mathfrak{I}}$ of $M$, together with a $T^{k_{i}}$-action on a finite normal (or Galois) cover $\pi_{i}\colon\tilde{U}_{i}\to U_{i}$ for each index, such that the torus actions satisfy certain compatibility conditions when passing from one neighborhood to another, and they are compatible with the action of the deck-transformation group of $\pi_{i}$ (see for example [20, Definition 19.2]). An $\mathsf{F}$-structure induces a partition of $M$ into submanifolds called the orbits of the $\mathsf{F}$-structure. We say that a sequence of Riemannian metrics $\{g_{n}\}_{n\in\mathbb{N}}$ on a compact manifold $M$ collapses with bounded curvature if there exist constants $\lambda\leq\Lambda\in\mathbb{R}$ such that
@@
and for any $p\in M$ the injectivity radius of $g_{n}$ at $p$ goes to 0 as $n\to\infty$ (see for example [41, p. 69]). By [6] and [7] we have that $\mathsf{F}$-structures characterize collapse with bounded curvature: A compact manifold collapses with bounded curvature if and only if it admits an $\mathsf{F}$-structure whose orbits have positive dimension.
For the proof of ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions, we prove that given a regular $B$-foliation $(M,\mathcal{F},g)$ on a compact manifold with leaves of positive dimension we may always collapse $M$ with bounded curvature and bounded diameter.
@@
Consider $(M,\mathcal{F},g)$ a regular non-trivial $A$-foliation on a compact Riemannian manifold, such that each leaf $L\in\mathcal{F}$ with the induced Riemannian metric $g|_{L}$ is flat, i.e. a $B$-foliation. Then by shrinking the directions tangent to the leaves of $\mathcal{F}$ we may collapse with bounded curvature and uniformly bounded diameter the manifold $(M,g)$.In particular we show there exists a family of Riemannian metrics $\{g_{\delta}\}_{1\geq\delta>0}$ on $M$ such that $\lambda\leq\mathrm{Sec}(g_{\delta})\leq\Lambda$, $\operatorname{Diam}(g_{\delta})\leq D$ and $\operatorname{Injrad}(g_{\delta})\to 0$ as $\delta\to 0$.
We recall that in the study of collapse with bounded curvature in [5], given such a collapsing sequence the authors construct $\mathsf{N}$-structures and highly symmetric $\mathsf{N}$-invariant Riemannian metrics which are bounded by the collapsing sequence (see ?THM? LABEL:T:_Bounded_collapse_in_1-connnected_manifolds). When the manifold $M$ has finite fundamental group an $\mathsf{N}$-structure is given by an $\mathsf{F}$-structure (see ?THM? LABEL:T:_N-structure_over_manifold_with_finite_fundamental_group_are_given_by_T-structure). In the case when $M$ has finite fundamental group we use the relationship between $\mathsf{N}$-structures and $\mathsf{F}$-structures, the invariant metrics and the explicit collapsing sequence given in the proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations to show that the geometry of $(M,\mathcal{F},g)$ where $\mathcal{F}$ is a regular $B$-foliation with leaves of positive dimension is approximated by geometries invariant under $\mathsf{F}$-structures. We can also point to a relationship between the tangent spaces of the leaves of the foliations and the tangent spaces of the orbits of the $\mathsf{F}$-structure.
@@
Consider $(M,\mathcal{F},g)$ a regular non trivial $A$-foliation on a compact $m$-dimensional Riemannian manifold with finite fundamental group and $m>2$, such that each leaf $L\in\mathcal{F}$ with the induced Riemannian metric $g|_{L}$ is flat, i.e. a $B$-foliation. Then there exists a sequence of Riemannian metrics $\{\bar{g}^{\varepsilon}_{N}(\varepsilon)\}_{1\geq\varepsilon>0}$ with each metric invariant under an $\mathsf{F}$-structure, and such that as $\varepsilon\to 0$, we have that $(M,\bar{g}^{\varepsilon}_{N}(\varepsilon))$ converges in the Gromov-Hausdorff sense to $(M,g)$ as $\varepsilon\to 0$. Moreover, given $p\in M$ fixed and taking $x\in T_{e}T^{k}$ fixed in the Lie algebra of the torus tracing the leaves of the $\mathsf{F}$-structure around $p$, we have that the action field $X^{\ast}(p)$ converges to a direction tangent to the leaf $L_{p}$.
Remark 1.1.
By ?THM? LABEL:L:_uniform_convergence_of_bar_g_to_g we also have that the Riemannian tensors $\bar{g}_{N(\varepsilon)}^{\varepsilon}$ converge uniformly to $g$.
We remark that by ?THM? LABEL:MT:_Collapse_of_flat_foliations, given a regular $B$-foliation on a compact manifold $M$ with leaves of positive dimension we have a continuous family $g_{\delta}$ of Riemannian metrics collapsing with bounded curvature to the leaf space. Combining this with [44, Theorem 0.4] we conclude that we have an upper bound for the minimum of the sectional curvature of a compact manifold equipped with a $B$-foliation whose leaves have positive dimension.
@@
Let $(M,\mathcal{F},g)$ be a regular $B$-foliation on a compact manifold with leaves of positive dimension. Then we have for $\lambda$ equal to the minimum of the sectional curvature of $g$, that $\lambda\leq 0$.
@@
?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions sheds some light on this question: Regular flat $A$-foliations (i.e. $B$-foliations) on simply-connected manifolds are given by torus actions, and thus are $\mathsf{F}$-structures.
We highlight that on a simply-connected manifold, an $\mathsf{F}$-structure is induced by a torus action (see for example [47, Proposition 3.1, Lemma 3.2 and Lemma 4.1]). Thus it is a pure $\mathsf{F}$-structure, and therefore a $B$-foliation. Still such $B$-foliations may have singular leaves. For example torus actions of cohomogeneity $2$ can have singular leaves [40], [8]. Nonetheless, when $M$ has finite fundamental group a regular $A$-foliation with flat leaves is the limit of $\mathsf{F}$-structures by ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures.
We point out that not all regular $A$-foliations admit a Riemannian metric making it a $B$-foliation. In [16] the authors construct manifolds of dimension $m\geq 9$ with non-trivial finite fundamental group and each equipped with a regular $A$-foliation whose leaves are homeomorphic to the (m4)(m-4)-dimensional torus but not diffeomorphic to the (m4)(m-4)-dimensional torus. These foliations do not admit any Riemannian metric making them into a $B$-foliation, since by [3] a flat manifold is homeomorphic to the torus if and only if it is diffeomorphic to the torus.
@@
Is there an $A$-foliation on a simply-connected manifold which is not given by a torus action?
@@
In [8, Theorem A] a negative answer to ?THM? LABEL:Q:_Question_2 is given for $A$-foliations of codimension $2$. Namely in [8], the author explored the problem of whether we can compare the diffeomorphism type of two compact simply-connected manifolds with $A$-foliations via their leaf spaces equipped with local information of the foliation, focusing on the case of $A$-foliations of codimension $2$. The positive answer given in [8, Theorem C and Lemma 5.11] to this comparison problem together with the classification of tours actions of cohomogeneity $2$ in [40], allows us to conclude that an $A$-foliation of codimension $2$ on a compact simply-connected manifold is given up to a foliated homeomorphism by a torus action. Then, by analyzing the local information of the foliation we are able to improve the conclusion to a foliated diffeomorphism.
@@
A crucial point in the proof of [8, Theorem A] is that for an $A$-foliation of codimension $2$ on a compact simply-connected manifold the stratification of the leaf space is simple and can be explicitly described. In general, even for singular Riemannian foliations on vector spaces with finite dimension, it is not clear how to compare the foliations via their leaf spaces (see [23, Question 1.2]). We point out that the proof presented in [8] is of a topological nature and only uses the fact that the presence of a foliated Riemannian metric guarantees a nice local structure of the foliation on tubular neighborhoods of leaves (see [35]). Moreover, to be able to extend the ideas in [8] to higher dimensions, we would require to impose extra topological conditions on $M$ and on the leaf space of the foliation. Thus, to the best of our knowledge, the approach presented in this manuscript is a different approach for comparing foliated structures to what exists in the literature.
@@
The present work was motivated by the need to give geometric conditions on $A$-foliations that force them to be given by torus actions. Due to [5] and [47] it follows that on simply-connected manifolds, torus actions with orbits of positive dimension are characterized by the phenomenon of collapse with bounded curvature (see ?THM? LABEL:T:_Bounded_collapse_in_1-connnected_manifolds). Thus, a natural approach to finding sufficient geometric conditions for an $A$-foliation to be induced by a group action, is to find sufficient geometric conditions which let us collapse with bounded curvature the manifold by shrinking the leaves of a given $A$-foliation on a simply-connected manifold.
@@
We highlight that to the best of our knowledge in the literature there are few references on the subject of deforming foliations. In [14] the authors show that for a particular class of $1$-dimensional regular foliations, a portion of the manifold has a cover into “long and thin” submanifolds extending the notions in [5] to the foliated setting. In [9] the authors consider a parametrized family of Lie groupoids and give rigidity results for such “deformations” of Lie groupoids. Lie groupoids are closely related to regular foliations (see [37]), and also singular Riemannian foliations (see [2]). Applying these rigidity results for Lie groupoids, the authors in [11] consider on a compact manifold a family of regular foliations $\mathcal{F}_{t}$ parametrized by a real parameter $t\in[0,1]$, each having compact leaves and Hausdorff leaf space, and show that at $t=0$ the existence of a leaf $L$ with $H^{1}(L,\mathbb{R})=0$ is a sufficient condition for such a parametrization to be rigid, i.e. $\mathcal{F}_{t}=\mathcal{F}_{0}$ for all $t\in[0,1]$. This is an analogous result to the conclusions presented in [13] and in [29]: Consider two regular foliations $\mathcal{F}_{1}$, and $\mathcal{F}_{2}$ with compact leaves on a fixed manifold $M$ without boundary, such that they are close to each other in an appropriate sense, and such that the leaf spaces of the foliations are Hausdorff spaces. In the case that there exists a leaf $L$ in one of the foliations with $H^{1}(L,\mathbb{R})=0$, there exists a foliated diffeomorphism between the foliated manifolds. For group actions there is a general rigidity result by Grove and Karcher [26]: Given a compact Lie group $G$, a compact manifold $M$, and two group actions $\mu_{1}\colon G\to\mathrm{Diff}(M)$ and $\mu_{2}\colon G\to\mathrm{Diff}(M)$, if the images $\mu_{1}(G)$ and $\mu_{2}(G)$ are $C^{1}$ close enough in $\mathrm{Diff}(M)$, then the images are conjugated in $\mathrm{Diff}(M)$.
@@
We point out that the proof of ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions is also a rigidity result for a deformation procedure, since the proof consist of approximating our foliation by torus actions using ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures, and showing that there exists a limit torus action whose orbits agree with the leaves of the foliation. Nonetheless, we point out that to the best of our knowledge there are no general rigidity results for singular Riemannian foliations as those presented in [11], [13], and [29]. Moreover, the rigidity result presented in ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions agree with the conclusions in [26]. One can also compare ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions with [44, Proposition 0.8] and comments after. In our work we also point to the existence of a fixed torus action on $M$, but we do show that we can take our foliated metric as the limit of the perturbed collapsing sequence.
@@
Our article is organized as follows: In Section 2 we review the necessary preliminaries on singular Riemannian foliations, as well as the concepts of $A$-foliations, $B$-foliations, and $\mathsf{N}$-structures. We also review Alexandrov spaces and equivariant convergence of metric spaces. At the end of Section 2, we present the relevant theorems of collapse theory used in our proofs. In Section 3 we give the proofs of our main theorems. We begin by presenting the proof of ?THM? LABEL:MT:_Collapse_of_flat_foliations, and then proceed with the proof of ?THM? LABEL:MT:_Approximating_flat_foliations_by_N-structures. We use then these results and their proofs to prove ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions.